A cashier has a total of 126 bills, made up of fives and tens. The total value of the money is $840. How many of each kind does he have?
Answer by stanbon(48569)

A cashier has a total of 126 bills, made up of fives and tens. The total value of the money is $840. How many of each kind does he have?
----------
Let # of $5 bills be x; These are worth $5x
# of 10 bills is 126-x; These are worth 10(126-x)=$(1260-10)
--------
EQUATION:
valur + value = 840
5x + 1260-10x=840
-5x=-420
x=84 (# of $5 bills)
126-x= 42 (# of $10 bills)
==========
Cheers,
Stan H.